# UniRGB-IR: A Unified Framework for Visible

### v1

@@ Abstract
Semantic analysis on visible (RGB) and infrared (IR) images has gained attention for its ability to be more accurate and robust under low-illumination and complex weather conditions. Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, existing methods prefer to design task-specific frameworks and directly fine-tune them with pre-trained foundation models on their RGB-IR semantic relevance datasets, which results in poor scalability and limited generalization. In this work, we propose a scalable and efficient framework called UniRGB-IR to unify RGB-IR downstream tasks, in which a novel adapter is developed to efficiently introduce richer RGB-IR features into the pre-trained RGB-based foundation model. Specifically, our framework consists of a vision transformer (ViT) foundation model, a Multi-modal Feature Pool (MFP) module and a Supplementary Feature Injector (SFI) module. The MFP and SFI modules cooperate with each other as an adpater to effectively complement the ViT features with the contextual multi-scale features. During training process, we freeze the entire foundation model to inherit prior knowledge and only optimize the MFP and SFI modules. Furthermore, to verify the effectiveness of our framework, we utilize the ViT-Base as the pre-trained foundation model to perform extensive experiments. Experimental results on various RGB-IR downstream tasks demonstrate that our method can achieve state-of-the-art performance. The source code and results are available at https://github.com/PoTsui99/UniRGB-IR.git
@@
RGB-IR Downstream Tasks, Multi-modal Fusion, Adapter, Transformer.
@@ I Introduction
Single visible (RGB) image semantic analysis is a common practice in computer vision and has been widely used in a variety of vision tasks [1, 2, 3]. With the development of physical devices, more and more general-purpose foundation backbones [4, 5] pre-trained on a large-scale dataset (e.g., ImageNet [6] and COCO [7]) have been designed for various downstream tasks. Since these pre-trained foundation models implicitly encode a large amount of prior knowledge, many researchers [8, 9] always fine-tune them on their semantic relevance datasets to significantly improve the performance on downstream tasks and speed up training convergence.
@@
However, visible cameras have proved to be difficult in providing reliable imaging [11, 12], due to their limited bandwidth across the spectrum. Especially in low-illumination and complex weather conditions, the useful information of RGB images is relatively scarce. Therefore, infrared (IR) modality with stronger low-light adaptability are used as additional information to supplement RGB modality [11, 13]. Afterwards, the joint use of RGB and IR images has been applied in more and more semantic analysis tasks.
@@
Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, most existing methods [14, 15] prevailing trend to use pre-trained RGB baselines and fine-tune them on their RGB-IR semantic relevance datasets as shown in Figure 1 (a). For example, RGB-IR object detectors [14, 16] utilize RGB-based models as strong baselines and fine-tune them to extract RGB and IR features. SwinNet [17] is explored based on the RGB-based transformer to extract hierarchical features of each modality for the detection of salient RGB-IR objects. Similarly, Ha et al. [18] propose a new baseline by aggregating infrared modality features into the RGB-based framework and retrain it in the RGB-IR semantic segmentation benchmark. Among the efforts to achieve competitive results, these approaches still suffer from two major issues. (1) Different downstream tasks require customized model structures, resulting in poor model scalability and excessive parameter storage. (2) Full fine-tuning strategy impairs the prior knowledge encoded in the foundation model, which limits the generalization potential of the fine-tuned model. Based on these issues, a question that naturally arises is whether we can adapt the RGB-based foundation model to efficiently construct a unified framework for RGB-IR downstream tasks.
@@
To achieve this purpose, in this paper, we explore an efficient and scalable framework for RGB-IR downstream tasks and, for the first time, propose a Unified model for RGB-IR downstream tasks called UniRGB-IR as shown in Figure 1 (b). Drawing inspiration from recent advances in adapters [10, 19], which is initially used to yield a extensible model to effectively exploit the potential representation of foundation models in natural language processing (NLP) field, we aim to develop an adapter to dynamically introduce richer RGB-IR features into the pre-trained foundation model through an additional CNN branch. Therefore, without modifying the original foundation model, the powerful pre-trained weights can be directly loaded and frozen to accelerate training convergence and achieve efficient fine-tuning on downstream tasks. Specifically, we design two core modules in our UniRGB-IR: a Multi-modal Feature Pool module (MFP) and a Supplementary Feature Injector module (SFI). In the MFP module, the multi-receptive field convolution and the feature pyramid are deployed to capture contextual multi-scale features from RGB and IR images. These features are fused at different scales to complement foundation models for different RGB-IR downstream tasks. For the SFI module, a cross-attention mechanism is used to dynamically inject the required features into the pre-trained foundation model, which facilitates our UniRGB-IR framework with a robust feature representation ability. To inherit prior knowledge of foundation model pre-trained on large-scale datasets, we utilize the adapter tuning paradigm instead of the full fine-tuning manner. In specific, we freeze the entire pre-trained weights and only optimize the MFP module and the SFI module. Therefore, our framework can be utilized as a unified framwork to efficiently fine-tune on various RGB-IR downstream tasks.
@@
  • We explore an efficient and scalable framework called UniRGB-IR to unify RGB-IR downstream tasks. To the best of our knowledge, this is the first attempt to construct a unified framework for RGB-IR downstream tasks.
  • We design a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter dynamically injects the required features into the pre-trained model. These two modules can be efficiently fine-tuned with adapter tuning paradigm to complement the pre-trained foundation model with richer RGB-IR features for RGB-IR downstream tasks.
@@ II-A Vision Foundation Models
Recently, thanks to the powerful long-distance modeling capability, vision transformers (ViT) [4] have been widely used as the foundation model in many vision tasks to achieve competitive results. The original ViT is a plain, non-hierarchical structure for image classification. Based on it, ViTDet [20] also constructs a non-hierarchical model by incorporating the feature pyramid. However, the non-hierarchical structure lacks rich feature representation, resulting in unsatisfactory performance. Subsequently, various hierarchical transformers [21, 22, 23, 5] have been proposed for different downstream vision tasks. Swin Transformer [21] and Multiscale Vision Transformer [22] are designed based on the ViT model to explore multi-scale features to improve the performance of image classification and object detection tasks. Besides, PVT [24] performs global attention on the downsampled key and value maps for dense prediction. In this paper, we also introduce the ViT model as the pre-trained foundation model to construct a unified framework for different RGB-IR downstream tasks.
II-B RGB-IR Downstream Tasks
Since infrared cameras have stronger low-light adaptability to capture object thermal radiation, infrared images have been widely used together with RGB images in various RGB-IR downstream tasks in recent years.
RGB-IR Object Detection. Zhang et al. [25] explore a two-stream SSD [26] structure to capture the contextual enhanced features for RGB-IR object detection. Besides, AR-CNN [27] is presented based on faster R-CNN [28] to align RGB and IR features. With the emergence of transformers, Yuan et al. [29] propose a complementary fusion transformer (CFT) module to achieve advanced detection results. Furthermore, C2Former [12] is a novel transformer block and can be incorporated into exited pre-trained models to increase intra- and inter-modality feature representations.
RGB-IR Semantic Segmentation. MFNet [18] is proposed to incorporate infrared features into the RGB-based framework to perform RGB-IR semantic segmentation. Based on the transformer structure, Wu et al. [30] propose a novel CCFFNet to excavate discriminative and complementary modality features for RGB-IR semantic segmentation. Moreover, CMX [31] is designed as a universal cross-modal fusion framework for RGB-IR semantic segmentation in an interactive fusion manner.
@@ II-B RGB-IR Downstream Tasks
The above methods attempt to design task-oriented structures to improve performance on corresponding downstream tasks. They either train the designed model from scratch or adopt a full fine-tuning strategy on a pre-trained model. Unlike these methods, we take a further step and deploy an adapter based on the pre-trained foundation model to unify the RGB-IR downstream tasks.
@@ III-B Multi-modal Feature Pool
Specifically, for the input RGB ($H\times W\times 3$) and IR ($H\times W$) images, we first employ a stem block borrowed from ResNet [41] to extract two modality features $F^{rgb}_{1}$ and $F^{ir}_{1}\in\mathbb{R}^{H/4\times W/4\times C}$. Then, these two features are split into four equal parts by utilizing the channel splitting technique [42]. To extract multi-receptive field perception, each part is subjected to convolution operations with different kernel sizes ($3\times 3$, $3\times 3$, $5\times 5$ and $7\times 7$). Then, we fuse each processed feature from two modalities by using a SE attention module (shown in Figure 3). Therefore, we concatenate each fused part to obtain RGB-IR contextual features $F_{fus}$. The above process can be represented as:
@@ III-D Adapter Tuning Paradigm
IV Experiments
To evaluate the effectiveness of our UniRGB-IR, we utilize the ViT-Base model (pre-trained on COCO [7] dataset) as the foundation model and utilize this framework to perform different RGB-IR downstream tasks. During training, we freeze the ViT-Base model and only optimize the MFP and SFI modules. We evaluate and compare our method with various competitive models, including CNN-based and Transformer-based models. Besides, our evaluation spans various tasks, including RGB-IR object detection on FLIR [70], LLVIP [71], and KAIST [53] datasets, RGB-IR semantic segmentation on MFNet [18] and PST9000 [63] datasets and RGB-IR salient object detection (see supplementary materials). Furthermore, ablation experiments on the designed modules and qualitative experiments are also conducted to verify that the UniRGB-IR framework can be leveraged as a unified framework to efficiently introduce IR features into the foundation model to achieve superior performance.
@@ IV-A RGB-IR Object Detection
Settings. All the experiments are conducted with NVIDIA GeForce RTX 3090 GPUs. We implement our framework on the MMDetection library and use the Cascade R-CNN [47] as the basic framework to perform RGB-IR object detection. The detector is trained with an initial learning rate of 2$\times$10-4 for 48 epochs. The batch size is set to 16, and the AdamW [72] optimizer is employed with a weight decay of 0.1. Horizontal flipping is also used for data augmentation.
@@
Results on KAIST dataset. The quantitative results of the different methods on the KAIST dataset are shown in Table II. The experiments are conducted under ‘All-dataset’ settings [53]. We compare our UniRGB-IR with thirteen multi-modal object detection methods. Our model achieves the best performance on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of five subsets (‘Near’, ‘Far’, ‘None’, ‘Partial’ and ‘Heavy’), and rank second in the ‘Medium’ subset. Furthermore, our detector surpasses the previous best competitor C2Former by 3.18% on the ‘All’ condition, which indicates UniRGB-IR is robust to the complex scenes. Intuitively, we draw the log-average Miss Rate (MR) over the False Positive Per Image (FPPI) curve of these detectors as shown in Figure 5. The above results verify the superiority of our UniRGB-IR.
@@
TABLE V: Quantitative comparison of RGB-IR salient object detection task on VT821, VT1000 and VT5000 datasets. * represents RGB-D SOD transformed into RGB-T SOD. The best results are highlighted in red and the second-best are highlighted in blue.

### v2

# UniRGB-IR: A Unified Framework for RGB-Infrared Semantic Tasks via Adapter Tuning

@@ Abstract
Semantic analysis on visible (RGB) and infrared (IR) images has gained attention for its ability to be more accurate and robust under low-illumination and complex weather conditions. Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, existing methods prefer to design task-specific frameworks and directly fine-tune them with pre-trained foundation models on their RGB-IR semantic relevance datasets, which results in poor scalability and limited generalization. In this work, we propose a general and efficient framework called UniRGB-IR to unify RGB-IR semantic tasks, in which a novel adapter is developed to efficiently introduce richer RGB-IR features into the pre-trained RGB-based foundation model. Specifically, our framework consists of a RGB-based foundation model, a Multi-modal Feature Pool (MFP) module and a Supplementary Feature Injector (SFI) module. The MFP and SFI modules cooperate with each other as an adapter to effectively complement the RGB-based features with the rich RGB-IR features. During training process, we freeze the entire foundation model to inherit prior knowledge and only optimize the proposed adapter. Furthermore, to verify the effectiveness of our framework, we utilize the vanilla vision transformer (ViT-Base) as the pre-trained foundation model to perform extensive experiments. Experimental results on various RGB-IR downstream tasks demonstrate that our method can achieve state-of-the-art performance. The source code and results are available at https://github.com/PoTsui99/UniRGB-IR.git
@@
RGB-IR Semantic Tasks, Multi-modal Fusion, Adapter, Vision Transformer.
@@ I Introduction
Image semantic analysis in the visible images is a common practice in computer vision and has been widely used in a variety of vision tasks [1, 2, 3]. However, visible cameras have proved to be difficult in providing reliable imaging [4, 5], due to their limited bandwidth across the spectrum. Particularly under low-illumination conditions and adverse weather scenarios, RGB images suffer from limited informative content. Therefore, infrared (IR) modality with stronger low-light adaptability is used as additional information to supplement RGB modality [4, 6]. Afterwards, the joint use of RGB and IR images has been applied in more and more semantic analysis tasks.
With the rapid advancements in physical computing capabilities, more and more general-purpose foundation backbones [7, 8] pre-trained on large-scale RGB-based datasets (e.g., ImageNet [9] and COCO [10]) have been designed for various tasks. Since these pre-trained base models implicitly encode a lot of prior knowledge, it is believed that pre-trained models can promote the improvement of downstream task performance and speed up training convergence. Therefore, fine-tuning the pre-trained foundation models [11, 12] on their semantic relevance datasets has become a common paradigm to improve the performance on specific downstream tasks.
@@
However, due to the lack of pre-trained foundation models on the large-scale infrared image datasets, for RGB-Infrared semantic tasks, a straightforward and prevailing way [14, 15] is to use pre-trained RGB-based models and fine-tune them on their RGB-IR semantic relevance datasets as shown in Figure 1 (a). For example, RGB-IR object detectors [14, 16] utilize RGB-based models as strong baselines and fine-tune them to extract RGB and IR features. SwinNet [17] is explored based on the RGB-based transformer to extract hierarchical features of each modality for the detection of salient RGB-IR objects. Similarly, Ha et al. [18] establish a novel baseline by incorporating infrared modality features into an RGB-based framework and fine-tune it on the RGB-IR semantic segmentation benchmark. Among the efforts to achieve competitive results, these approaches still suffer from two major issues:
(1) Different semantic tasks require customized model structures, resulting in poor model versatility and excessive parameter storage.
(2) Full fine-tuning strategy impairs the prior knowledge encoded in the foundation model, which limits the generalization potential of the fine-tuned foundation model.
Above challenges necessitate the adaptation of RGB-based foundation models to construct a unified framework capable of handling RGB-IR semantic tasks effectively.
@@
Drawing inspiration from recent advances in adapters [13, 19], which are initially used to yield an extensible model to effectively exploit the potential representation of foundation models in the natural language processing (NLP) field, we develop an adapter to dynamically introduce richer RGB-IR features into the pre-trained RGB-based foundation model. In our adapter, a feature extractor is designed to obtain rich RGB-IR features and a feature injector is proposed to adaptively introduce the required features for the foundation model. Without modifying the original RGB-based foundation model, the powerful pre-trained weights can be directly frozen to accelerate training convergence and achieve efficient fine-tuning on downstream tasks. Therefore, in this paper, we propose an efficient and universal foundation model for RGB-IR semantic tasks. In particular, we are the first to construct a Unified framework for RGB-IR semantic tasks, called UniRGB-IR, as shown in Figure 1 (b).
Specifically, we design two core modules in our UniRGB-IR: a Multi-modal Feature Pool module (MFP) and a Supplementary Feature Injector module (SFI). In the MFP module, the multi-receptive field convolution and the feature pyramid are deployed to capture contextual multi-scale features from RGB and IR images. These features are fused at different scales to complement foundation models for different RGB-IR semantic tasks. For the SFI module, a cross-attention mechanism is used to dynamically inject the required features into the pre-trained foundation model, which facilitates our UniRGB-IR framework with a robust feature representation ability. To inherit prior knowledge of the foundation model pre-trained on large-scale datasets, we utilize the adapter tuning paradigm instead of the full fine-tuning manner. We freeze the entire pre-trained weights and only optimize our proposed adapter. Consequently, UniRGB-IR can serve as a unified framework to achieve effective fine-tuning for various RGB-IR semantic tasks.
@@
  • We explore an efficient and universal framework called UniRGB-IR for RGB-IR semantic tasks. To the best of our knowledge, this is the first attempt to construct a unified framework for RGB-IR downstream tasks.
  • We design a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter dynamically injects the required features into the pre-trained model. These two modules can be efficiently fine-tuned with adapter tuning paradigm to complement the pre-trained foundation model with richer RGB-IR features for RGB-IR semantic tasks.
@@
The remainder of the article is organized as follows. In the Section II, we review the scientific literature related to our proposed approach. Section III introduces our proposed UniRGB-IR for RGB-Infrared semantic tasks. Experimental results are presented in Section IV, along with the comparison with state-of-the-art methods. Finally, we conclude with a discussion of our contribution in Section VI.
@@ II-A Vision Foundation Models
Recently, thanks to the powerful long-distance modeling capability, vision transformers (ViT) [7] have been widely used as the foundation model in many vision tasks to achieve competitive results. The original ViT is a plain, non-hierarchical structure for image classification. Based on it, ViTDet [20] also constructs a non-hierarchical model by incorporating the feature pyramid. However, the non-hierarchical structure lacks rich feature representation, resulting in unsatisfactory performance. Subsequently, various hierarchical transformers [21, 22, 23, 8] have been proposed for different downstream vision tasks. Swin Transformer [21] and Multiscale Vision Transformer [22] are designed based on the ViT model to explore multi-scale features to improve the performance of image classification and object detection tasks. Besides, PVT [24] performs global attention on the downsampled key and value maps for dense prediction. In this paper, we also introduce the ViT model as the pre-trained foundation model to construct a unified framework for RGB-IR semantic tasks.
II-B RGB-IR Semantic Tasks
RGB-IR Object Detection. Zhang et al. [25] explore a two-stream SSD [26] structure to capture the contextual enhanced features for RGB-IR object detection. Besides, AR-CNN [27] is presented based on faster R-CNN [28] to align RGB and IR features. With the emergence of transformers, Yuan et al. [29] propose a complementary fusion transformer (CFT) module to achieve advanced detection results. Furthermore, C2Former [5] is a novel transformer block and can be incorporated into existing pre-trained models to increase intra- and inter-modality feature representations.
RGB-IR Semantic Segmentation. MFNet [18] is proposed to incorporate infrared features into the RGB-based framework to perform RGB-IR semantic segmentation. Based on the transformer structure, Wu et al. [30] propose a novel CCFFNet to excavate complementary modality features for RGB-IR semantic segmentation. Moreover, CMX [31] is designed as a universal cross-modal fusion framework for RGB-IR semantic segmentation in an interactive fusion manner.
@@ II-B RGB-IR Semantic Tasks
The above methods attempt to design task-oriented structures to improve performance on corresponding downstream tasks. They either train the designed model from scratch or adopt a full fine-tuning strategy on a pre-trained model. Unlike these methods, we take a further step and deploy an adapter based on the pre-trained foundation model to unify the RGB-IR semantic tasks.
@@ III-B Multi-modal Feature Pool
Specifically, for the input RGB ($H\times W\times 3$) and IR ($H\times W$) images, we first employ a stem block borrowed from ResNet to extract two modality features $F^{rgb}_{1}$ and $F^{ir}_{1}\in\mathbb{R}^{H/4\times W/4\times C}$. Then, these two features are split into four equal parts by utilizing the channel splitting technique [41]. To extract multi-receptive field perception, each part is subjected to convolution operations with different kernel sizes ($3\times 3$, $3\times 3$, $5\times 5$ and $7\times 7$). Then, we fuse each processed feature from two modalities by using a SE attention module (shown in Figure 3). Therefore, we concatenate each fused part to obtain RGB-IR contextual features $F_{fus}$. The above process can be represented as:
@@ III-D Adapter Tuning Paradigm
IV Object Detection and Semantic Segmentation
To evaluate the effectiveness of our UniRGB-IR, we utilize the ViT-Base model (pre-trained on COCO [10] dataset) as the foundation model and utilize this framework to perform different RGB-IR semantic tasks. During training, we freeze the ViT-Base model and only optimize the MFP and SFI modules. We evaluate and compare our method with various competitive models, including CNN-based and Transformer-based models. Besides, our evaluation spans various tasks, including RGB-IR object detection on FLIR [69], LLVIP [70], and KAIST [52] datasets, RGB-IR semantic segmentation on MFNet [18] and PST9000 [62] datasets. Furthermore, ablation experiments on the designed modules and qualitative experiments are also conducted to verify that the UniRGB-IR framework can be leveraged as a unified framework to efficiently introduce IR features into the foundation model to achieve superior performance.
@@ IV-A RGB-IR Object Detection
Settings. All the experiments are conducted with NVIDIA GeForce RTX 3090 GPUs. We implement our framework on the MMDetection library and use the Cascade R-CNN as the basic framework to perform RGB-IR object detection. The detector is trained with an initial learning rate of 2$\times$10-4 for 48 epochs. The batch size is set to 16, and the AdamW optimizer is employed with a weight decay of 0.1. Horizontal flipping is also used for data augmentation.
@@
Results on KAIST dataset. The quantitative results of the different methods on the KAIST dataset are shown in Table II. The experiments are conducted under ‘All-dataset’ settings [52]. We compare our UniRGB-IR with thirteen multi-modal object detection methods. Our model achieves the best performance on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of five subsets (‘Near’, ‘Far’, ‘None’, ‘Partial’ and ‘Heavy’), and rank second in the ‘Medium’ subset. Furthermore, our detector surpasses the previous best competitor C2Former by 3.18% on the ‘All’ condition, which indicates UniRGB-IR is robust to the complex scenes. Intuitively, we draw the log-average Miss Rate (MR) over the False Positive Per Image (FPPI) curve of these detectors as shown in Figure 5. All these results verify that our UniRGB-IR can exploit the potential representation from the RGB-based foundation model to achieve superior results.
@@
TABLE V: Ablation studies of key components on the FLIR and PST900 datasets. The best results are highlighted in bold.
FLIR PST900
MFP SFI mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$ mAcc mIoU
39.6 76.3 35.9 90.1 77.5
41.2 77.9 36.5 91.7 78.8
44.1 81.4 40.2 94.3 82.8
IV-B RGB-IR Semantic Segmentation
Results. The quantitative results of the different RGB-IR segmentation methods on the PST900 and MFNet dataset are shown in Table III and Table IV respectively. The comparison results show that our model significantly outperforms other methods. Specifically, our model obtains the best performance in terms of both the mACC and mIoU. In PST900 dataset, our model performs competitive performance in the Backpack, Hand-Drill and Survivor, outperforming the second-best methods by 2.3%, 0.6% and 0.4% IoU, respectively. In predicting Color Cone class on MFNet dataset, our model achieves the first and second best in terms of both evaluation metrics, with about 11% Acc higher than that of the second-best. These results show that our UniRGB-IR can also be used flexibly for segmentation tasks without any modification, which validates the universality of our method.
TABLE VI: Ablation of adding SFI module to different stages.
Different Stages with SFI module mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$
Stage 1 Stage 2 Stage 3 Stage 4
41.7 80.0 37.2
43.3 81.6 39.7
44.1 81.4 40.2
42.7 79.9 38.2
TABLE VII: Ablation of the different attention mechanisms.
Attention Mechanism Complexity mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$
Global Attention [82] Quadratic 36.1 72.8 30.6
Pale Attention [42] Linear 31.6 68.4 24.6
Deformable Attention [43] Linear 44.1 81.4 40.2
IV-C Ablation Study
Ablation for components. To investigate the contribution of the SFI and MFP modules, we gradually add each module to the baseline model as shown in Table V. We stack RGB and IR images and feed them into the Cascade R-CNN detector and SETR framework (ViT backbone) for full fine-tuning on the corresponding tasks, which are considered as our baseline models. Then, we freeze the ViT model and introduce the MFP module to extract contextual multi-scale features (element-wise addition), which results in 1.6% mAP and 1.3% mIoU improvements. Finally, we replace the element-wise addition operation with the SFI module and further improve the mAP and mIoU metrics by 2.9% and 4.0% respectively, which achieve the best performance on both datasets.
SFI module at different stages. We add the SFI module to the ViT pre-trained model at the beginning of different stages. From Table VI, we can find that by adding the SFI module at the first stage, the detector achieves 41.7% mAP on FLIR dataset. After adding the SFI module in the second and third stages, the performance further improved by about 2% and 3% mAP, respectively. However, continuing to add it to the final stage will reduce the detection performance while also increasing computational overhead. Therefore, we add the SFI module from the first stage to the third stage of ViT model.
Attention type in SFI module. Since the attention mechanism in our SFI module is replaceable, we adopt three popular attention mechanisms in our UniRGB-IR to discuss their impact on model performance. As shown in Table VII, the detector achieves the best performance with linear complexity by utilizing deformable attention. Thus, the deformable attention is more suitable for our framework and utilized as the default configuration. It is worth noting that it can be replaced by other attention mechanisms to further achieve superior performance.
Figure 7: Training efficiency analysis on the FLIR dataset.
IV-D Visualization Analysis
[…]
TABLE VIII: Quantitative comparison of RGB-IR salient object detection task on VT821, VT1000 and VT5000 datasets. * represents RGB-D SOD transformed into RGB-T SOD. The best results are highlighted in red and the second-best are highlighted in blue.